Thirteen decreased by three times a number is sixteen. I am not sure how to work this one. I have tried what I know how to do, but i amd confused.
Answer by checkley77(12844)

13-3X=16
-3X=16-13
-3X=3
X=3/-3
X=-1 ANS.
PROOF:
13-3*-1=16
13=3=16
16=16